UT May 14 2022 - Ful DEC Night



Night Info

  • Sunset 19:10
  • Nautical Twilight Ends 20:10
  • Astro Twilight Ends 20:44
  • Night Center 00:16
  • Astro Twilight Starts 03:47
  • Nautical Twilight Starts 04:22
  • Moon sets 04:26
  • Sunrise 05:21

Staff

  • OSA: S.Allanson
  • ISp:
  • Mtn:
  • ISA: B.Rothberg, R.T.Gatto
  • AO: J.Power
  • Special Guests: LBTI team

Night Plan

  • Starting configuration LBTI
  • LBTI DX Com/Eng Checkout - until tasks complete - High Priority
  • Test leaving DX Config in place for ASM - Verify no impact to SL observations with flat - High Priority
  • Reconfigure for bino-LUCI, authorize LUCI-AO/ESM
    • LUCI-Bino regardless of conditions
  • LUCI-1 AO/ESM
    • OSCO - 0.5hr
    • ESM observations of faint Bin 4 targets - 0.5-0.75hr
      • Already Generated Mixed Mode Scripts: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/Bin4_ESM_Faint_Test
      • DONE GD 321 - R=16.6 WFSMag=17.7
        • GD321_Bin4_R16pt6_ESM_SL.xml - mixed mode; GD321_Bin4_R16pt6_L1.xml - mono
        • twiligh-03:15hr
        • 9pt dither H-band - 1min shutter time per dither
      • DONE SBSS 1510+510 R=17.1 WFSMag=17.96
        • GD321_Bin4_R16pt6_ESM_SL.xml - mixed mode; SBSS1510+510_Bin4_R17pt1_L1.xml - mono
        • visible all night
        • 9pt dither H-band - 1min shutter time per dither
      • DONE 6C 165641+474241 QSO Rmag=17 (18.2)
        • 6C165641_QSO_Bin4_R17_ESM_SL.xml - mixed mode; 6C165641_QSO_Bin4_R17_L1.xml - mono
        • visible all night
        • 9pt dither H-band - 1min shutter time per dither
    • ESM Characterization (if time) - 1hr
      • DONE M57 Ring Nebula - 23:30-twilight - FeII and PaBeta
      • M92 - 21hr-twilight
    • Angular Separation tests - 0.5-1hr
      • in OT: AO STARS - FULL DITHERS *NEW*-->Binaries
      • May only do some
        • Target 2 BD+00 2888 R=10.7 - separation 0.225" - need 256x2048 subframe - visible twilight-00:30hr
          • Target 5 GD 135 R=14.6 - separation 0.3" - no subframe - visible twilight-23:30hr
        • Target 4 SBSS 1148+564 R=14.9 separation 0.6" - no subframe - visible twilight - 02hr
        • Target 6 2MASSJ1128+3136 R=13.5 - 1" separation - no subframe - visible twilight - 01hr
        • Target 7 2MASSJ13061537+2043444 R=8.5 - 1.6" separation -64x2048 subframe - visible twilight-02:15hr
          • Note: given separation and lack of GS only 1 PA script is available - companion should be visible directly below main star
          • Target 9 2MASSJ10590395+6349283 Bin 2 R=12.7 - 1.57" separation - No subframe backup - visible twilight-01:30hr
        • Target 3 BD+33 2137 R=10.3 - 2.4" separation no subframe
          • Target 10 2MASSJ15553957+4025135 Bin2/3 R=13.3 - 2.3" separation - visible 01:15-twlight
    • DONE D-time Science AO/ESM - Engineering Sub-frame test
      • mid-z ULIRG FF1725
        • ESM Bin 1 AO Ref=7.4 - SL in parallel - 0.75hr Ks
        • seeing < 1.2"

Poor Weather Backup

  • LUCI-Bino (or mono)
  • Rothberg Local Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V47.xml
      • 1 hour per target - 1hr15 with telluric
      • Not in order of priority - observe best placed target
      • NGC 4004 - DEC +27 - twilght-02hr
      • UGC 9829 - DEC -01 - 21:30-twilight
  • Photometric Standards - any seeing, but must be clear
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
    • Select stars based on visibility, need to get stars at two different airmasses


Night Log

Closed Dome

  • UT 00:27 - starting LUCI GUIs, init all, test darks
    • L1 0001-0002 - 2.51sec dark
    • L2 0001-0002 - 2.51sec dark
  • UT 00:29 - moving N30Fieldstop into place for LUCI1
  • UT 00:32 - moving N30camera into place
  • UT 00:33 - checking fieldstop
    • L1 0003-0005
  • UT 00:35 - ready for LUCI-1 AO
  • UT 00:36 - checking LUCI-2 N375 fieldstop
    • L2 0003-0006
  • UT 00:40 - ready for LUCI-2 SL N375

On-Sky

LBTI
  • UT 02:06 - opening dome
  • Starting with LBTI Eng
  • UT 07:53 - SX ASM fail
  • UT 07:55 - LBTI handing over to LBTO
LUCI AO/ESM
  • UT 08:03 - putting N30Fieldstop away, S.Allanson moving to pointing star and collimation star
Bin 4 ESM Faintness Test

GD321
  • UT 08:16 - sending preset Bino
  • Rmag = 16.7, WFSmag = 17.6
  • L1 3.54pix = 0.42" , L2 5.52 pix = 0.66"
  • Frame L1 0009 - 0.46"
  • Frame L1 0012 - 0.39" L2 0013 - 0.43" - have to do a more thorough check
    • L1 0006-0013
    • L2 0007-0015

SBSS1510+510
  • UT 08:36 - sending preset Bino
  • Rmag= 17.1, WFSmag = 18.0
  • UT 08:43 - loops closed
  • Frame 0014 - L1 ~0.38", Frame 0015 L2 - ~0.46" - however, ellipticity L1 0.12, L2 0.46
    • L1 0014-0021
    • L2 0015-0022

6C 165641+474241 - this is a QSO - on-axis Bin4 of a 17th mag QSO!
  • UT 08:57 - sending preset
  • Rmag = 17, WFSMag=18.3
  • UT 09:04 - loops closed
  • First pair, L1 = 0.42", L2 = 0.51"
  • on-souce 2.4 photons/subapp, 1.5photons/subapp for sky
  • taking TN for L1 0025
  • taking TN for L1 0029
    • L1 0022-0029
    • L2 0023-0030
M57 Ring Nebula - ESM characterization
  • UT 09:19 - sending preset
  • Rmag = 12.2, WFSMag = 14.2
  • sent as Bin1, but its Bin 2 - required manual intervention, auto would have failed
  • UT 09:32 - loops closed
  • First exposure L1 ~0.38", L2 ~0.45" ellip L1 0.04, L2 0.11 - Guider on DX 0.49"
  • Frames L1 0036, L2 0037, 0.38" vs 0.43", DX Guider shows 0.7"
  • Frames L1 0038, L2 0039, 0.49" vs 0.53" , DIMM 1.3", DX Guider shows 0.71"
  • SEEING BUBBLE (from DIMM)
  • UT 10:05 - DIMM 1.5, DX Guider 0.9-1"
  • Pa Beta (J) - AO Ref star - with worse seeing L1 0.58", L2 0.87"
  • UT 10:11 DX Guider 1.3"
  • L1 #0052 0.69", L2#0053 0.78"
  • L1 #0056 0.56", L2#0057 0.93"
    • L1 0030-0059
    • L2 0031-0060
  • UT 10:25 - done
FF1725 Bin 1/Sub-frame Test
  • Rmag = 7.3, WFSMag = 8.9 (Kmag=5.5)
  • UT 10:27 - sending preset
  • Do sub-frame 1024x2048 readout to avoid K=5.5 AO Ref star
  • Back down to 1" on DIMM and 0.9" on DX Guider
  • UT 10:37- loops closed
    • L1 0060-0061 - acq
    • L2 0061-0062 - acq
  • Subwin 1035,1036, 1004,1039 - 1024x2048
    • L1 0062-0064
    • L2 0063-0065
  • swtich to 768x2048 - sent to GEIRS without re-drawing - itime is now 0.94 DIT x 60 NDIT
  • L1 0.43", L2 0.66"
    • L1 0065 -0070
    • L2 0066-0071
  • switch to 512x2048 - 0.62 DIT x 60 NDIT - 60"x4' - dither is +/-45" random
    • seems to be working - don't see the crosstalk - L1 FWHM ~0.44"
  • UT 11:08 - DX guider 0.8"
  • L1 Frame 75-74 - FWHM ~0.38"
  • L1 Frame 81-80 FWHM ~0.44"
  • L1 frame 85-84 FWHM~0.44"
  • L1 frame 88-87 FWHM ~0.42"
    • L1 0071-0094
    • L2 0072-0095
    • Last frame L1 ~0.41"FWHM
  • Sub-frame to mask out works
  • UT 11:35 - stopping & closing

Calibrations Closed Dome

  • N375 Imaging Flats
    • PaBeta
      • L1 0095-0104
      • L2 0096-0105
    • FeII
      • L1 0105-0114
      • L2 0106-0115
    • H-band
      • L1 0115-0124
      • L2 0116-0125
    • Ks-band
      • L1 0125-0134
      • L2 0126-0135
  • UT 12:05 - putting both LUCIs in safe config - ALL DONE

Summary

  • Handover to LUCI-ESM testing ~01hr local time
  • Pushed the limits of the faint end of Bin 4 to WFSMag=18.3 (closed loop on 17.6 and 18.0 as well)
    • 20 min per object - no issues in large offsets and returns to field or small dithers
    • IQ FWHM~0.38"-0.5" (further processing will be done to verify) - in some cases the same IQ was achieved with Full AO and ESM Bin 4 for the same target
    • One target was a QSO - achieved and maintained on-axis observations of WFSMag=18.3 QSO for ~20min without issue
  • Continued characterization of ESM using Ring Nebula field
  • Tested using sub-frame readout in field with AO Ref star K=5.5 (WFSMag=8.9)
    • sub-framed around target to avoid impact of bright star saturating star (using integration times 0.6-1sec)
    • diffraction spikes and reflections did not seem to impact observations
    • changed sub-frame size "on-the-fly" - Paused, decreased sub-frame, continued with no impact
    • observations continued past 12deg twilight with no change in IQ at Ks-band, IQ stable across observations 0.38"-0.41"
    • tested simultaneous sub-frame readout (60 NDITS both LUCIs in parallel)

-- BarryRothberg - 13

May 2022
Topic revision: r10 - 14 May 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback